The product of twenty less than eight times a number and thirty less than two times a number equals zero. Find the number(s).
Answer by mananth(16949)

(8x-20)(2x-30)=0
16x^2-240x-40x+600=0
16x^2-280x +600 =0
divide by 8
2x^2-35x+75=0
2x^2-30x-5x+75=0
2x(x-15)-5(x-15)=0
(x-15)(2x-5)=0
x=15 OR 5/2